### v1

# Clinicians’ Voice: Fundamental Considerations for XAI in Healthcare

Abstract
Explainable AI (XAI) holds the promise of advancing the implementation and adoption of AI-based tools in practice, especially in high-stakes environments like healthcare. However, most of the current research is disconnected from its practical applications and lacks input of end users. To address this, we conducted semi-structured interviews with clinicians to discuss their thoughts, hopes, and concerns. We find that clinicians generally think positively about developing AI-based tools for clinical practice, but they have concerns about how these will fit into their workflow and how it will impact clinician-patient relations. We further identify education of clinicians on AI as a crucial factor for the success of AI in healthcare and highlight aspects clinicians are looking for in (X)AI-based tools. In contrast to other studies, we take on a holistic and exploratory perspective to identify general requirements, which is necessary before moving on to testing specific (X)AI products for healthcare.
Keywords explainable AI, trustworthy AI, human-centered explainable AI, healthcare, interview study
1 INTRODUCTION
Rapid developments in artificial intelligence (AI) call for more transparency of AI-based technologies to tackle concerns related to trustworthiness, responsibility, and accountability. This has led to the emergence of the field of Explainable AI (XAI). The goal of XAI is making AI models more transparent and understandable to address these concerns and aid the implementation of AI-based technologies. Especially in high-stakes situations, like those in healthcare, AI-based systems could enhance critical decision-making processes. Ideally, AI models provide a certain prediction (e.g., the probability of a skin lesion being cancerous) which supports the clinicians in their decision process. However, beyond regulatory requirements these technologies have to meet, there is an apparent reluctance to use AI-based tools as it is unclear how the AI model arrives at its prediction. Although user studies, evaluations, and multidisciplinary research are generally advocated for within XAI, only a small fraction of publications connect their work with practical applications. With this research, we contribute to the research on XAI in healthcare by conducting semi-structured interviews with clinicians from the Netherlands. These interviews are aimed at developing a better understanding of the current state of (X)AI in healthcare, and clinicians’ perception of the development of AI-based tools. We identify clinicians’ hopes and concerns for (X)AI in healthcare and formulate concrete takeaways to inform research and development in the field. Aspects we looked at include desirable characteristics of XAI methods/explanations, tasks that clinicians (would) find (X)AI helpful for, concerns and fears they face regarding the implementation of (X)AI in their workflow, and their perception of the role of trust in the implementation and adoption of (X)AI. We confirm findings from previous research, including that users emphasize the importance of receiving information about the background of the model in addition to its predictions and explanations. Our results further highlight the importance of multidisciplinary projects to ensure seamless introduction of (X)AI-based tools into clinical practice, and the significance of educating clinicians on AI before implementing AI-based tools in their workflow.
2 BACKGROUND
One can generally distinguish between glass-box and black-box models in XAI. Glass-box models are transparent, and model internals and operations are accessible to the users such that they can trace back every prediction made by the model. Thus, glass-box models provide both local and global explainability. Black-box models are opaque machine learning models where the user is unable to trace back the prediction due to the model complexity. Typically, (only) local explainability is achieved by complementing single predictions made by black-box models with post-hoc explanations. For an in-depth account of developments in XAI, we refer to several recent review papers (see, for example, Das et al., 2020; Gilpin et al., 2018; Murdoch et al., 2019; Ras et al., 2022; Linardatos et al., 2021; Guidotti et al., 2018; Carvalho et al., 2019; Barredo Arrieta et al., 2020; Ali et al., 2023)
In healthcare, critical decisions are made daily and hence, the potential for AI-based decision support systems is undeniable. This is also reflected in the rising number of publications on AI-based clinical decision support tools (DSTs) that incorporate explainability. For example, Thoral et al. (2021) trained an AI-model to predict re-admission or death within seven days after discharge from the ICU and explain predictions using feature importance scores Lundberg and Lee (2017). Both Abdullah et al. (2021) and Antoniadi et al. (2021) provide reviews of explainable machine learning approaches in healthcare including examples of publications on medical AI tools with explainability. Markus et al. (2021) focus on how explainability can be evaluated and provide a step-by-step guide to choose the most appropriate XAI method. Amann et al. (2020) take a multidisciplinary approach to determining the need for explainability in healthcare from technological, legal, medical, patient, and ethical perspectives. They conclude that there is an apparent need for explainibility and amongst others highlight the importance of considering the interaction between healthcare professionals and AI-based DSTs.
Cai et al. (2019) conduct a user study including semi-structured interviews on AI-assisted diagnosis of prostate cancer, focusing on identifying desired aspects for explainability in the AI tool. They find that pathologists are interested in global properties of the AI tool beyond accuracy, such as how conservative or liberal its decisions are or what kind of data it was trained on. In a similar fashion, Naiseh et al. (2023) empirically evaluated the effect of different types of explanations on trust calibration in a medical context and conducted interviews with some participants after the experiment. Like Cai et al. (2019), they find that practitioners are interested in global information of AI-tools, such as the contents of the data and the recency of the trained model. They further find that practitioners worry about fitting AI-tools into their workflow because it adds extra workload and that they would like more guidance on how to interpret the explanations. Tonekaboni et al. (2019) conducted interviews with 10 clinicians that are familiar with AI-based clinical tools, primarily to identify what explanation types best serve their needs and to propose evaluation metrics of explainability tools. For example, they found that clinicians value feature importance scores as they draw attention to specific patient characteristics relevant for the prediction. They further conclude that example-based explanations are less helpful as they depend on the definition of similarity with the added difficulty that seemingly similar patients may have significantly different outcomes and vice versa. They also find that accuracy alone is not sufficient and that the output should be complemented with a certainty score for the prediction.
Other studies in the realm of XAI in healthcare have a slightly different focus. For example, Lee et al. (2024) address onboarding material for XAI-tools, Subramanian et al. (2024) focus on a broader stakeholder perspective not only involving clinicians, and Bienefeld et al. (2023) analyse collaboration between developers and clinicians. Our study differs from this previous research in several ways. First, we do not build our work around a specific case or AI-tool. In doing so, we eliminate the factor of expertise from the study design. Second, we interview clinicians from different fields within healthcare to record general data. Third, our study is exploratory in nature rather than empirically driven. That is, we do not test for an effect of XAI methods on, for example, trust or understanding. The goal of our research is to identify clinicians’ requirements for successful implementation of AI-tools in clinical practice with the aim of improving decision making in healthcare.
3 METHODOLOGY
We conducted 11 semi-structured interviews with open-ended questions with clinicians from the Netherlands practicisng in various fields. The study was approved by the ethical body of our institution and all participants gave their informed consent prior to participation. An interview scheme with prompts and guiding questions was used as basis for the interviews (see Table 1). All interviews were conducted by the same interviewer to limit threats to validity and reliability introduced by interviewer variability. Upon completion of the interviews, each interview was transcribed and then analysed using a coding scheme. We did not pre-define any coding categories as this is an exploratory study. Instead, we collected relevant quotes from the transcribed interviews for each of the topics addressed and then identified common themes. A more detailed account of our methodology including participants and recruitment, procedure, materials, analysis, and limitations is given in Appendix A.
Table 1: Interview scheme
Topic Questions
Opening
(a) Establishing rapport [interviewer thanks participant for their time; interviewer introduces themselves]
(b) Purpose of the study [interviewer shortly introduces XAI and explains purpose of the study]
(c) Procedure [interview will last about 30 minutes, an audio recording will be started]
Body
(a) Background Could you tell me a bit more about field of work and your expertise?
What specialization do you work in?
How long have you been working in this field?
(b) AI in healthcare In healthcare generally, and also in your field, are AI systems in use?
Could you describe the extent to which you are confronted with automatic systems that use AI in your daily work?
For what type of tasks are AI systems used?
Can you tell me more about your research on AI for […]?
Can you think of AI systems that are not (yet) in use but that you could imagine helpful?
What do you believe are the reasons why implementation of AI in practice is slow?
(c) Attitude towards AI Could you describe your positive and negative feelings about using AI in healthcare practice?
What are your thoughts on trust in AI?
(d) Explanation methods [interviewer gives overview and example of explanation methods using a toy example]
For each of these methods, can you briefly describe how familiar you are with them?
What are your thoughts on each of these methods?
What are desirable characteristics of an explanation for an AI system?
Closing
(a) Final comments Is there anything else that you would like to share with me on this topic?
Do you have any questions to me?
(b) Closing [interviewer thanks interviewee again for their time and participation]
4 RESULTS
4.1 Current use of AI in healthcare
Based on the interviews, we can group AI-based tools into those related to (1) business operations, (2) care processes, and (3) medical device regulations. The first category of AI systems is relatively universal to any larger business, not only hospitals or care facilities, and covers aspects like finance, forecasting, and so on. The second category relates to AI systems that enhance the care process by, for instance, making it more efficient. A typical example that several interviewees provided was the planning of chemotherapy. The last category of AI systems in healthcare concerns AI systems that fall under the medical device regulation. Simply put, a medical device is any instrument intended to be used to diagnose, treat, monitor or prevent diseases “and does not achieve its primary intended action by pharmacological, immunological or metabolic means” (International Medical Device Regulators Forum (2013), IMDRF, p.7). AI models fall under the category Software as a Medical Device, which includes software that is intended to be used for any of the aforementioned purposes without being part of a hardware medical device (International Medical Device Regulators Forum (2013), IMDRF). In layman’s terms, decision support tool is more commonly used to describe AI as a medical device. Here, an AI model is trained to make a prediction based on patient data and hence, supports clinicians in their decisions.
When asked to what extent AI models are implemented in medical care in their field specifically, interviewees unanimously said that there is very little AI (as DSTs) integrated in their workflows. Most interviewees pointed out that a significant amount of budget is invested in research and development of medical AI models. Though many of these efforts are being tested, very few of them actually reach the bedside. Arguably, the most popular field for AI applications within healthcare is radiology, with the number of publications on AI in radiology being “sky high” (interview 4). However, even though there are more than 200 products in AI for radiology available on the European market [^1] https://radiology.healthairegister.com, not many of those are integrated in clinical practice yet. Furthermore, one interviewee pointed out that most of these products perform specific subtasks, such as segmentation or quantification tasks, instead of predicting a concrete medical outcome; thus, only indirectly supporting clinicians.
However, some interviewees also mentioned that other forms of AI models, such as those for people planning and scheduling, are integrated everywhere, and considering AI in that sense “everybody is really already using AI a lot”, but “not […] as a medical device yet” (interview 10). Note that this falls into the second category of AI models previously mentioned.
4.2 What tasks should AI be used for?
Even though some interviewees held the opinion that AI for administrative tasks is already widely used, others said that they would want to see AI be more used for administrative and logistics tasks. One interviewee said that “80% of our task is administration” (interview 6), hence if AI could be used for this type of work, clinicians could focus more on their patients. Two interviewees mentioned a type of large language model for medical purposes, for example to help them write summary reports which is a tedious task taking up a significant amount of time.
Regarding AI as a clinical DST, four interviewees mentioned predicting complications as a useful application. For example, when doing surgery, they would like to know upfront how big the chance is to run into a certain complication. And even beyond that, one interviewee said they “would like to know upfront what can go wrong and how can we prevent that” (interview 5). Hence, AI could serve not only as a warning system but also as a guide on how to navigate certain situations. To give a concrete example, one interviewee from obstetrics and gynaecology explained that when a woman is in labor, they monitor both the mother and the baby, meaning they collect a large dataset. Having an AI model trained on this data that continuously predicts the chance of a birth ending in a C-section would save them not only time and resources, but would also reduce the risks for women in labor. Another example comes from oncology, where it would be beneficial to predict the likelihood of complications during surgery but also the probability of remission following treatment.
Two interviewees mentioned that monitoring or medical surveillance would be helpful, especially in cases where patients are not continuously at the hospital. With the recent technological advancements, it may be possible to provide a wide range of patients with wearables to monitor their health. In this case, (X)AI-based tools could be used to predict time of admission to the hospital or make recommendations on further treatment, which could then be evaluated in consultations and check-ups with the clinician.
4.3 Attitude towards (X)AI in healthcare
Positive attitude.
The most common response was that AI is simply more efficient than humans and can “make our life easier” (interview 3). Again, it was highlighted that AI can take over largely the tedious tasks, hence acting not necessarily as a DST but rather taking over relatively well-defined but time-consuming tasks, such as planning and writing reports.
All interviewees had mainly positive thoughts about AI-based DSTs, most of them believing that AI can help them make the best decision. As a human being, “it’s really difficult to make complex decisions and to take into account more than four or five variables” (interview 9), which is also supported by research in cognitive sciences (e.g., Miller, 1956; Cowan, 2001; Luck and Vogel, 1997; Oberauer and Kliegl, 2006). Working memory capacity has also been linked to decision-making, and several works have shown that humans tend to rely on heuristics to make decisions if cognitive demands are too high (e.g., Bröder and Schiffer, 2003; Shah and Oppenheimer, 2008). However, it is not only about how many variables clinicians need to take into account when assessing the patient’s situation, but also relating these to their domain knowledge and previous cases. For example, one interviewee pointed out that AI “has a memory of a lot more cases than I can experience and remember” (interview 2), so it can base its prediction on a much wider range of previous patients. Another was convinced that “you can treat a patient better with the data you have” (interview 5), which will “make it better” (interview 5) for both the patient and the doctor.
While interviewees seemed to agree that AI can support them in the process of making the best decision for their patients, two explicitly mentioned that AI could do so by taking a more objective stance. That is, sometimes our judgments can be clouded by emotions and especially in fields where there is frequent clinician-patient contact one may have too much of a “tunnel vision on [a] patient” (interview 3), in which cases it would be helpful to have “an objective somebody next to you” (interview 3).
Negative attitude.
However, interviewees also voiced concerns that the distance between them and their patients will become larger when using AI-based DSTs. The same case may have different outcomes because of the patients’ wishes, beliefs, and preferences – “soft features” (interview 2) that are not captured in the data. Personalised treatment is only possible if you interact and connect to your patient, know their story, and incorporate their wishes. Using AI-based devices for some tasks bears the fear that the distance between doctors and their patients becomes so large that they will be dependent on these devices. This outcome would take away a big part of what drives a lot of clinicians, which is simply the interaction with their patients. However, it is important to note that other interviewees, when asked about their opinion on this aspect, said that they are not concerned about it because regardless of where any medical information comes from, it is still them (the clinicians) who are in charge of the situation and in control of the decision. As the name suggests, a DST is merely there to give support – the decision still lies with the clinician who takes them in consultation with the patient.
Bias and AI literacy are other aspects that concern clinicians. In the context of AI, bias relates to the data the model was trained on, and has been an important topic in recent research, especially after criticisms were voiced against some algorithms implemented in practice; see, for example, BBC (2018). Bias in AI exists when the model favors certain people, or provides more accurate predictions for a certain group of individuals. In healthcare, this could mean that an AI-based DST may provide more accurate guidance for a certain group of patients. For example, if the dataset used for training included a higher proportion of old people, the model might give misleading or incorrect results for younger patients. If such systems were to be used, clinicians should be aware of these drawbacks to accurately interpret the output given by the device. Several interviewees pointed this out and referred to it as validity. In fact, five interviewees mentioned that the validation of AI-based devices is extremely important and difficult, making it hard to develop models that can be applied to all patients. Overall, interviewees in this study showed that they are aware of the issue, and further pointed out that this is especially concerning if clinicians rely too much on the AI model instead of using it complementary to their expert knowledge. This concern was voiced especially for younger generations of clinicians who are and will be developing their expertise while these systems are around, potentially making them less critical towards them.
Generally, participants seem to agree that AI is the future and can solve many problems. However, is this really true? To quote one of the interviewees, AI is “kind of the Holy Grail to a lot of people” (interview 6), without considering the context. In a significant number of cases, expert knowledge and/or simple models may be just as good as complex AI models, with the added benefit of providing concrete insights into the decision-making process Semenova et al. (2022), which is also something several interviewees acknowledged.
4.4 What’s best – feature importances, counterfactual explanations, or rule-based methods?
Comparing only feature importance scores, counterfactual explanations and rule-based models, there is a clear tendency for people to prefer feature importance scores. Interviewees generally agreed that it is intuitive, quick and easy to understand, as you can grasp it all with one glance. It is so easy that “you don’t need to be a professional to understand what’s going on here” (interview 3). While this is clearly a pro of this method, it also raises concerns – “the […] question is, is it really then that simple?” (interview 5). Especially in complex decisions that require consideration of a multitude of factors, it is almost “not believable” (interview 1). Interviewees also pointed out that they are familiar with this method of explaining the output of AI models, and hence may be biased in preferring it over the other two methods.
Counterfactual explanations require a bit more time to be understood, however they can provide one with more in-depth insights. As one interviewee put it, this type of explanation is more “fluid” (interview 5), rather than static as the feature importance scores, and it shows the effect of something changing. This property could also be used to improve prognosis, which however raises the question of causality – a change in features leads to a different prediction by the model, which does not always translate to reality. Lastly, one interviewer explained that looking at counterfactuals means one is essentially looking at other patients, and not focusing on the current patient anymore. On top of that, there are endless other patients out there, so why would one look at this specific counterfactual?
Rule-based methods are similar to how decisions are made –e.g., “Of course, if you are a two year old, you will never have a mama carcinoma”, or breast cancer in layman’s terms (interview 3)–, and they’re helpful if you really want to understand the prediction. Knowing the decision mechanism makes the AI’s prediction more “believable” (interview 1), gives a sense of control, and something concrete to base your decision on, which is especially helpful in high stakes decisions. One interviewee held the opinion that such simple model should always be preferred and we should only move towards complex AI models (and with that other types of explanations) if the simple models are not enough. However, two interviewees also pointed out that if you have many rules, this may be too complicated to use on the spot. For research purposes, when you want to uncover the mechanism behind patient characteristics and medical conditions or outcomes, rule-based methods are helpful – but in clinical practice?
Regardless of the type of explanation, visual explanations are preferred for being easily understood on the spot, and because it’s a way to remove any potential language barriers. Another aspect interviewees raised was uncertainty of the model prediction – being shown how certain the model is about a prediction makes it more trustworthy and helps clinicians navigate the usage of the information given. This ties in with knowing more about the background of the model, such as which data were used for training and testing, how it was validated, how big the sample was, and so on, which came up in about half of the interviews.
4.5 The role of trust in AI
One third of the interviewees mentioned trust comes automatically when a device is tested, certified, and approved. This is comparable to trusting lab results and equipment without exactly knowing how they work, or when they have last been calibrated. Somewhat in line with that, trust develops over time as you use a device and experience its benefits. Some people are hesitant towards new products, known in economics and sociology as late majority and laggards Rogers (2003). Especially older clinicians seem to be reluctant to use new AI-based medical devices. One interviewee raised the point that trust is distinct from responsibility, that is you can trust a device but in the end you have to take responsibility for your decisions, and “how are you going to feel okay with that responsibility when it’s not your own thought?” (interview 3).
5 DISCUSSION
Based on the results obtained, we highlight the opportunities for (X)AI research, the requirements for successful implementation of (X)AI in practice, and the main warnings provided regarding (X)AI usage and development.
5.1 Opportunities for (X)AI research:
Applications in administrative tasks, risk estimation and warning systems, and out-of-hospital health monitoring
: The interviews have shown a positive attitude towards further (X)AI development for clinical practice. Application directions that are mentioned more frequently are lowering the administrative burden for medical personnel, forecasting complications and estimating potential risks and risk prevention, early warning systems, and tools that help monitor patients while they are out-of-hospital, for example via wearable devices.
Development of visual explanations for AI models:
In terms of ease-of-use, visual explanations are indicated to be relatively easy to interpret as they remove potential language barriers. This means there is potential for research focused on the development of visual explanations in XAI.
5.2 Requirements for successful implementation of (X)AI in practice:
More emphasis on multidisciplinary projects is needed for the development of (X)AI techniques:
While there is a large body of research on (X)AI-based tools for clinical practice, only very few make it to the bedside. Many reasons for this slow implementation can be attributed to a mismatch between all the different parties involved. Multidisciplinary collaboration is key to building a successful product, and Van de Sande et al. (2022) propose a structured approach to and discuss steps that are needed to develop and safely implement AI in medicine.
(X)AI training is required for clinicians:
Most interviewees in this study have some knowledge about (X)AI, however, mainly for their own interest in the topic. As it stands momentarily, medical education in the Netherlands does not include obligatory classes on AI or AI-based medical devices, which can have significant consequences. First, understanding the background of the model, such as the training data used and how it was validated, is important to judge its performance and get a sense of trust. Second, some explanations may be too simplistic and giving a false sense of understanding, like feature importance plots. It is important for the user to understand what explanations really show and how to interpret them. Third, it is often mistaken that explanations imply causality. Using AI-based tools without sufficient background knowledge makes the users prone to misinterpret the information provided.
Besides, without proper training, we might run the risk of two groups forming among clinicians: those who readily accept and trust AI-based tools, and those who are too resistant to use it. The first group might simply not be aware of important background information that could impact the usage and interpretation of the output. The second group might be too skeptical and hesitant to use technology they do not understand. Both could end up in effects opposite to what we want to achieve with the implementation of AI-based DSTs in clinical practice. Hence, education is key to a successful implementation of these tools as acknowledged in previous work Tjoa and Guan (2021).
(X)AI tools should include more than just explanations:
It is apparent that clinicians desire more than a simple explanation of the prediction, as many have pointed out that they would like to see a certainty score, which is also in line with findings of previous work Tonekaboni et al. (2019). Besides, we can confirm findings of previous work Cai et al. (2019); Lee et al. (2024) that suggest that clinicians are interested in the background of the model, such as information about the training data or benchmark performances. Transparency on training data and representativeness are also important to build trust in the model. A comparison can be made with the setup of clinical trials and evidence-based medicine.
We need more awareness of change management and how to navigate it:
The world is changing at a high pace and especially technological progress is faster than ever, but humans are known to resist change and adapt slowly. From a neurological perspective, humans have been shown to respond faster to familiar stimuli Yang and Beck (2023) and hence tend to stick to droutines. In the workplace, this means that we tend to resist changing our routine or workflow that we are used to. A lot of barriers in the adoption of AI-based DSTs are psychological in nature, as one of the interviewee says (interview 8), and hence we need to draw on knowledge from that field to overcome these barriers. Hence, the development of products for clinical practice needs to be linked to research on human cognition and psychology to allow for a fast and smooth adoption of the technology.
However, while we should make sure to incorporate research from other fields into the development of these products to aid their implementation, people also need to experience its practical benefits over time and need to get used to using them. By now, no-one questions the lab equipment, whereas there likely was some skepticism when they were first introduced.
5.3 Warnings regarding (X)AI usage and development:
Black-box models should not be used if they are not a significant improvement over glass-box models:
AI is supposed to be the future, but is it a good idea to use it everywhere? Lately, the field receives more and more criticism for being blindly applied to any project, just because it is AI and it looks good. However, there are many situations where a much simpler model can achieve the same level of performance as a complex black-box model Semenova et al. (2022). Especially when it comes to high stakes situations, one should first use an interpretable glass-box model which allows clinicians to understand the mechanism, and only resort to more complex AI models if they sufficiently outperform glass-box models.
AI should not increase the distance between clinicians and patients:
The goal of AI-based DSTs is to enhance medical care, and not to act as middleman between clinicians and patients. Hence, if there is a disconnect between both parties due to the introduction of such a technology, then we have missed the goal. DSTs in clinical practice should be intuitive to use to the extent that you should fail to notice their existence in the workflow. This has been coined as ‘unremarkableness’ by Yang et al. (2019), who studied the degree of unremarkableness of AI to fit into the clinical decision-making process. To achieve this, we need contextual awareness early on in the design and development process.
6 CONCLUSION
This study offers insights from practicing clinicians on the integration of (X)AI in healthcare. Through semi-structured interviews, clinicians highlighted the potential benefits and challenges of (X)AI in clinical settings. While AI can improve decision-making, concerns were raised about the reliance on complex (X)AI models, which may erode trust and create distance between clinicians and patients. It is essential to strike a balance between AI’s capabilities and human judgment to uphold patient-centered care.
One crucial takeaway is the importance of seamlessly integrating (X)AI into clinical workflows to enhance decision-making without causing disruptions. Clinicians emphasized the need for AI literacy and stressed the significance of collaboration among AI developers, healthcare professionals, and clinicians for successful adoption.
The study also delved into clinician preferences for explanation methods, with feature importance scores being favored for their simplicity and ease of comprehension. However, counterfactual explanations and rule-based methods were valued for providing deeper insights in intricate cases. Visual explanations were deemed essential by clinicians as they enhance the intuitiveness and accessibility of AI outputs, particularly in time-sensitive situations. The choice of explanation method should align with the clinical context — feature importances may be suitable for quick decisions, while decision rules or counterfactual explanations are more appropriate for detailed analysis or audits.
In conclusion, this study advocates for transparent (X)AI systems that prioritize visual clarity, multidisciplinary collaboration, thoughtful selection of explanation methods, and integration as supportive tools rather than barriers in healthcare practice based on real-world clinician input.
While this study offers valuable insights into clinicians’ perspectives on XAI in healthcare, it has several limitations, including its narrow geographical focus on the Netherlands, reliance on qualitative data, potential selection bias, and lack of patient input or practical implementation. We believe this study provides a strong groundwork for future research addressing these limitations.
Appendix A METHODOLOGY
We conducted 11 semi-structured interviews with open-ended questions with clinicians trained and working in the Netherlands. In this section we discuss recruitment strategy, interview procedure, material and limitations of our design. The supplementary materials for the interviews are available in the appendix. The study was approved by the ethical body of our institution [to be filled in more detailed after review to preserve this as an anonymous submission]. Interviewees were informed of the goal and procedure of the interviews and signed an informed consent form prior to their participation.
A.1 Participants and recruitment
The main inclusion criterion for participants was that they had successfully completed their medical study in the Netherlands. The sample included clinicians from different fields within medicine (general practice, (pediatric) oncology, radiology, general health psychology, obstetrics and gynaecology, surgery) and with years of experience in their fields varying from two to 30 years. Participants were recruited through networking and word of mouth. None of the interviewees was personally known to the interviewer before the interview was conducted.
A.2 Procedure
Potential interviewees were contacted via email, which stated the reason for reaching out, the person that had provided the contact details, practical information such as the expected time, and the informed consent form. If interviewees responded positively, a follow-up exchange followed to arrange a meeting. Interviews were scheduled for 30 minutes, and the majority of them took place online as participants come from across the Netherlands. Participants were asked to sign the informed consent form prior to starting the interview. During the interview, the interviewer confirmed with participants that they agree to having the interview audio-recorded. Interviewees could ask questions or withdraw from the interview at any moment, and they were explicitly given the opportunity to add any information they wished to mention or ask questions at the end of the interview. All interviews were conducted by the same interviewer to avoid potential differences due to interviewing style.
A.3 Materials
An interview scheme with guiding questions was developed to provide structure and guidance to the interviews (see Table 1). At the start of each interview, the interviewer provided context for the interview after introductions, similar to what participants had read in the informed consent form beforehand. Then, the first set of questions was aimed at inquiring about the interviewee’s background, including the field they specialized in and experience in the field. The following questions were aimed at understanding the current use of AI in their field and healthcare in general. Besides, most interviewees are involved in research that oftentimes includes research on AI for a specific task in their field, and we asked them to share more about that, or other research they’re aware of. Then, we also asked them to come up with use cases of AI in their field which they may find helpful in their job. We also asked what they think are the reasons why these do not yet exist, and why implementation of AI in healthcare in general is slow. The next set of questions was focused on uncovering interviewee’s attitude towards AI in healthcare. This entailed asking specifically for their positive and negative feelings towards using AI in healthcare, and if they could describe their hopes and concerns. In this part of the interview, we also asked them about the role they believe trust plays for the implementation and adoption of (X)AI in practice. Then, for the next set of questions, the interviewer explained three different types of explanations using a simple toy example for illustration (see Appendix B) and asked interviewees to explain their thoughts considering each of these explanations. The explanations that we chose were (1) feature importance scores, (2) counterfactual explanations, and (3) a glass box model. The first method was chosen as feature importance is arguably the most popular method within XAI momentarily. Counterfactual explanations have received considerable attention lately, with many papers proposing algorithms to generate these types of explanations (see, for example, Russell, 2019; Ustun et al., 2019; Kanamori et al., 2021a, b; Karimi et al., 2020, 2021; Mothilal et al., 2020; Mahajan et al., 2019; Maragno et al., 2022). Research from the social and cognitive sciences shows that counterfactual explanations mimic the way humans give explanations themselves, and hence are expected to be good explanations in XAI (Byrne, 2019; Miller, 2019). Lastly, as a contrast, we opted for a glass-box model (weighted rules), such as those generated by Röber et al. (2024), as the third type of explanation. Besides their opinion on these specific explanation types, we included a question on what characteristics (e.g., visual, interactive) they would like to see in an AI system, or what type of information they would like to be shown. Then, at the end of the interview, interviewees were asked if they would like to share any other considerations with us.
A.4 Analysis
Upon the completion of all interviews, each one was transcribed and coded by the interviewer. This involved creating a coding scheme based on the interview scheme and the topics covered in most interviews. The coding scheme was then populated with relevant quotes and excerpts from the transcriptions to allow for a nuanced understanding of the interviewees’ responses. As this was an exploratory study, no predefined coding categories were used; instead, responses were carefully analyzed to identify recurring themes and patterns and the codes were developed organically in response to the data.
A.5 Limitations
The nature of our study bears some limitations we should consider when interpreting the results. First, interviews are known to bear threats to validity and reliability Alshenqeeti (2014). As a response to this, we chose to have the same interviewer conduct all interviews. Interviews were further audio-recorded and then transcribed for the analysis, by that avoiding misinformation through note-taking or incorrect understanding by the interviewer. Furthermore, during the interviews, time was taken to clarify any unclarities to avoid misunderstanding and misinterpretation of the responses. Second, as an explorative interview study, the sample size is limited and we did not perform any quantitative analyses of the results and should be careful with making generalizations. Third, the interviews were scheduled to last 30 minutes, which may not be considered long enough to discuss different aspects of XAI. As this is an exploratory study and clinicians are known to have limited time, we were constrained by this duration. Lastly, we reached out to candidates via email, and in some cases a common connection made the introduction. Generally, it can be expected that individuals interested in XAI are more likely to positively respond to an interview request. Hence, we may only have interviewed individuals with strong (positive or negative) opinions on the topic, by that missing out on the broader group of clinicians.
Appendix B INTERVIEW ADDITIONAL MATERIAL
The material below has been used during interviews to illustrate and explain three types of XAI methods: feature importances, counterfactual explanations, and rule-based methods.
Types of Explanations for Machine Learning
In machine learning, we distinguish between glass-box models vs. black-box models. Glass-box models are those that are inherently interpretable, i.e., we see the model and are able to grasp how it is making predictions. Black-box models are complex machine learning models that we are not able to interpret directly. For those models, we see the predictions and then apply methods that generate post-hoc explanations. In this study we are focusing on the most prevalent types of explanations used: weighted rules as glass-box model, and feature importances and counterfactual explanations as post-hoc explanations for black-box models.
Figure 1: Overview of XAI methods
Toy example
For the sake of exemplifying the type of explanations, consider this toy example: We have a dataset consisting of 200 intensive care patients. For each patient, different medical parameters such as respiratory rate, oxygen saturation, systolic blood pressure, and so on, have been measured. We call such parameters features and they’re named X1 to X5 in the dataset for simplicity. Please note that the values are intentionally chosen arbitrarily to keep the discussion general and avoid expert knowledge to factor in. These features are used to predict a binary (0 or 1) target variable like mortality, i.e., not-deceased (0) or deceased (1). A machine learning model has been fit to the data to predict the binary target variable. An excerpt of the data is shown below where the last column corresponds to the prediction of the model for the target variable. The highlighted row is the patient of interest, i.e., the patient for which the prediction is to be explained.
Figure 2: Excerpt of toy example data
Types of explanations
Feature importances
We see here the contribution of each feature to the average prediction visualised in a bar plot, for the highlighted patient in the example dataset. The prediction for that patient is 0 (not-deceased), and from the plot we can see that X5 contributed most to this prediction. The colors of the bars, red and blue respectively, indicate whether a contribution is positive (towards predicting 1) or negative (towards predicting 0). In this example, a positive contribution means that it is a risk increasing factor for death in this model, and a negative contribution indicates a risk decreasing factor. The length of the bar corresponds to the weight of the contribution. In this example, X5 has the biggest impact and is a risk decreasing factor.
Figure 3: Feature importances (SHAP) example
Counterfactual explanations
Counterfactual explanations show how the feature values of a patient should change such that the prediction would also change. The instance of which you want to explain the prediction is called the factual. In our case, the factual is the highlighted patient in the toy example, for whom the prediction is 0 (not-deceased). A counterfactual is another (possibly fictitious) patient, who is very similar to the factual with regard to the feature values (say equal heart rate, systolic blood pressure, and respiratory rate, but slightly different temperature and oxygen saturation), who is however predicted as 1 (deceased).
In the table below, this is visualised. The factual is the highlighted patient from our dataset that we want to derive an explanation for. The counterfactual is another patient, with equal values for X2, X4, and X5 as compared to the factual. However, for this other patient, X1 has a value of 2 and X3 has a value of 5. The prediction for this patient is 1 (deceased).
Figure 4: Counterfactual explanation example
Weighted rules
Instead of explaining a prediction made by a machine learning model, we can derive predictions ourselves using a set of rules. These rules follow the form: “If … AND …, then PREDICT … .”. For example,“IF systolic blood pressure is higher than 130 AND respiratory rate is higher than 25, then predict 1 (deceased”. If a rule does not apply to a patient, i.e., the IF-conditions are not met, this rule is not used for the patient. To predict the outcome for a patient, we use all rules for which the patient fulfils the IF-conditions. Each rule makes a prediction, which can be contradicting each other – some rules may predict 0 (not-deceased), some may predict 1 (deceased). Each rule comes with a weight, which can be thought of as the importance of the rules. For each class (not- deceased vs. deceased), we sum the weights of the respective rules. The final prediction will be the class with the largest sum of weights.
For the toy example the algorithm returns four rules, visually displayed here:
Figure 5: Weighted rules example
Using these rules, we can make the prediction for the highlighted patient: According to rule (1) and (2), the prediction for the highlighted patient would be class 0 (not-deceased). The summed weight for these rules is 1.0 + 0.3 = 1.3. According to rule (3), it would be predicted as class 1 (deceased), with a weight of 0.7. Rule (4) is not used for our patient, since the first condition of the rule (i.e., X1 < 3) is not met by the patient who has a value of X1 = 4. Hence, we use only the first three rules. To make the final prediction, we look at the summed weights: 1.3 and 0.7 for class 0 (not-deceased) and 1 (deceased), respectively. Hence the final prediction for the patient is 0 (not- deceased).